OpenStudy (anonymous):

25.2 mL of water are placed in a graduated cylinder. A 22.6 g stone is dropped in, and the water level rises to 32.4 mL. Find the density.

OpenStudy (anonymous):

Initial Volume = V1 = 25.2 mL ; Final volume = V2 = 32.4 mL Increase in volume = V2 - V1 = 32.4 mL - 25.2 mL = 7.2 mL as u know 1 mL = 1 cm^3 so density = mass / Volume = 22.6 g / 7.2 cm^3 = 3.138888889 g/cm^3

OpenStudy (anonymous):

@livelaughswim7